### v2

# FlowBind: Efficient Any-to-Any Generation with Bidirectional Flows

Abstract
Any-to-any generation seeks to translate between arbitrary subsets of modalities, enabling flexible cross-modal synthesis. Despite recent success, existing flow-based approaches are challenged by their inefficiency, as they require large-scale datasets often with restrictive pairing constraints, incur high computational cost from modeling joint distribution, and rely on complex multi-stage training. We propose FlowBind, an efficient framework for any-to-any generation. Our approach is distinguished by its simplicity: it learns a shared latent space capturing cross-modal information, with modality-specific invertible flows bridging this latent to each modality. Both components are optimized jointly under a single flow-matching objective, and at inference the invertible flows act as encoders and decoders for direct translation across modalities. By factorizing interactions through the shared latent, FlowBind naturally leverages arbitrary subsets of modalities for training, and achieves competitive generation quality while substantially reducing data requirements and computational cost. Experiments on text, image, and audio demonstrate that FlowBind attains comparable quality while requiring up to 6× fewer parameters and training 10× faster than prior methods. The project page with code is available at https://yeonwoo378.github.io/official_flowbind.
1 Introduction
Recent progress in flow-based generative models has delivered state-of-the-art performance in multi-modal generation. By conditioning on a given input modality, these models excel at specialist tasks such as text-to-image esser2024sd3; BlackForestLabs2024FLUX1 or text-to-audio synthesis liu2024audioldm2; huang2023makeanaudio2, demonstrating their strength in learning continuous cross-modal transformations. However, these successes are largely confined to fixed input and output mapping, and extending flow models to support true any-to-any generation, where arbitrary subsets of modalities can be generated given any other subsets, remains an open challenge.
Bridging the gap from specialist to generalist flow models introduces fundamental hurdles, primarily due to requirements of multi-modal data and computational cost. Frameworks that rely on a central anchor modality, typically text (tang2023codi), require each modality to be paired with text during training so that all modalities can be aligned through the shared text representation. This design is restrictive, as it prevents the model from learning the rich, direct correlations that exist beyond language. Conversely, methods that model the full joint conditioning of all modalities (li2025omniflow) can achieve expressive generation performance but at a steep cost: they require some fully-paired data for stable training, which is scarce, and their computational complexity often scales quadratically with the number of modalities. These data and compute issues render them impractical for real-world scenarios with a large and diverse set of modalities.
Beyond the computational cost, a significant hurdle for generalist models is the complexity of their training pipelines. Rather than a single, unified process, these frameworks often rely on intricate, multi-stage procedures. These stages separately optimize the encoding components for modality alignment and the decoding components responsible for the model’s generative capabilities. This staged approach is evident in prominent models; for instance, CoDi (tang2023codi) employs a multi-stage process that separates modality alignment from joint generation. Similarly, OmniFlow (li2025omniflow) requires a distinct post-training phase after merging its core components. Such multi-stage pipelines can be brittle, difficult to optimize, and hinder the development of truly seamless, end-to-end generative models.
We introduce FlowBind, a simple flow-based model that addresses these limitations. FlowBind introduces a learnable shared latent capturing cross-modal commonality, and connects each modality to this latent through its own invertible flow. All components are trained jointly under a single flow matching objective, while the learned flows enable direct any-to-any translation at inference. Because each flow requires only its modality paired with the latent, the method naturally supports training with partially paired data while reducing computational cost. This design yields a simple, efficient, and data-flexible solution for general-purpose any-to-any generation.
In summary, our main contributions are as follows: (1) We introduce a flow-based framework for any-to-any generation that factorizes multi-modal interactions through a learnable shared latent, enabling training from arbitrary paired data with low computation budget. (2) Our method jointly optimizes both the shared latent and all modality-specific flows under a single flow-matching loss, avoiding the multi-stage pipelines. (3) Experiments on text, image, and audio demonstrate that FlowBind achieves competitive quality with substantially reduced data and computation compared to representative baselines, while flexibly supporting any-to-any translation.
2 Preliminaries
Flow Matching
Conditional Flow Matching (lipman2023fm) is a simulation-free framework for learning a continuous transformation between a source distribution $p_{0}$ and a target distribution $p_{1}$. This transformation is defined by an Ordinary Differential Equation (ODE), $\frac{dz_{t}}{dt}=v_{\theta}(z_{t},t)$, where a drift network $v_{\theta}$ parametrizes the time-dependent vector field. With linear interpolation path $z_{t}\;=\;(1-t)\,z_{0}+t\,z_{1}$ with $(z_{0},z_{1})\sim(p_{0},p_{1})$, the target velocity is simply $z_{1}-z_{0}$, and the objective becomes:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}\left[\left\|v_{\theta}(z_{t},t)-(z_{1}-z_{0})\right\|^{2}\right].$ (1)
At the optimum, Eq. 1 yields the conditional expectation of the target velocity:
$v_{\theta}^{\star}(x,t)\;=\;\mathbb{E}\!\left[\,z_{1}-z_{0}\;\middle|\;z_{t}=x\;\right].$ (2)
Generation is then performed by integrating the learned drifts over time:
$z_{t_{1}}=z_{t_{0}}+\int_{t_{0}}^{t_{1}}v_{\theta}(z_{t},t)\,dt=\text{ODESolve}(z_{t_{0}},v_{\theta},t_{0},t_{1})$ (3)
Note that, under standard Lipschitz conditions, the induced flow is invertible i.e., the ODE can be integrated forward or backward in time to induce samples from $p_{0}$ or $p_{1}$.
Any-to-Any Generative Flows
The goal of any-to-any generation is to learn a unified model that can translate between arbitrary subsets of modalities. Given $N$ modalities $\mathbf{z}=(z^{1},\dots,z^{N})$, this amounts to modeling their joint distribution $p(\mathbf{z})$ so that for any $S_{\text{in}},S_{\text{out}}\subseteq\{1,\dots,N\}$, the model can perform any-to-any generation by sampling from conditional probability $p(\mathbf{z}^{S_{\text{out}}}|\mathbf{z}^{S_{\text{in}}})$.
Existing flow-based approaches address this problem by constructing continuous trajectories that transform i.i.d. Gaussian noise $\mathbf{z}_{0}\sim\pi_{\text{prior}}$ into data samples $\mathbf{z}_{1}\sim\pi_{\text{data}}$. Representative examples include CoDi tang2023codi and OmniFlow li2025omniflow, which mainly differ in how they synchronize trajectories across modalities. CoDi learns modality-specific encoders that align all modalities to a shared text embedding, which then serves as the conditioning signal for per-modality denoising networks $\epsilon^{i}(z^{i}_{t},t,\hat{c}^{\text{text}})$. In contrast, OmniFlow learns a time-decoupled joint velocity field $v(z^{1}_{t_{1}},\dots,z^{N}_{t_{N}},t_{1},\dots,t_{N})$, where the interpolation path for each modality is explicitly conditioned with the other modalities to ensure alignment.
Despite their empirical success, existing flow-based methods face several limitations. First, they cannot fully leverage arbitrary paired modalities for any-to-any generation: CoDi requires each modality to be paired with text to establish a canonical embedding, while OmniFlow relies heavily on fully-paired data (i.e., samples where all $N$ modalities are present) for stable training  [^1] Although partially paired data can be used for training in principle, performance and stability are reported to depend strongly on fully paired data; see Appendix B.2 of li2025omniflow.. Second, both methods require multi-stage training: CoDi separately learns the shared representation and denoising networks, whereas OmniFlow pre-trains drift networks for each modality pair before joint training. Finally, they operate in high-dimensional representations, leading to substantial computational cost and slow convergence.
3 FlowBind
To address the aforementioned challenges, we propose FlowBind, a unified flow-based framework for any-to-any generation. FlowBind is designed to overcome key drawbacks of prior methods: it supports single unified training procedure, operates with lower computational overhead, and fully exploits partially paired data for effective learning.
The key idea of FlowBind is to replace the fixed Gaussian prior with a learnable, shared distribution that encapsulates common information across modalities. This acts as a latent anchor, where each modality is connected to it via their own invertible, per-modality flows (Figure 1). With this factorization, FlowBind achieves alignment across modalities naturally via the shared distribution, unlike existing approaches that anchor all modalities to text tang2023codi or couples them through a joint drift li2025omniflow. Meanwhile, both the shared distribution and the per-modality flows are learned jointly with only standard flow matching loss using partially paired data.
Formally, consider a subset of multi-modal data $\mathbf{z}^{S}=\{z^{i}|i\in S\}$ with $S\subset\{1,\dots,N\}$, which is sampled from a joint distribution $\mathbf{z}^{S}\sim\pi^{S}_{\text{data}}$. Assume that there exists a shared latent $z^{*}\sim\pi^{S}_{\text{shared}}$ that encompasses the common information of all individual modality in $\mathbf{z}^{S}$. Then for each $i\in S$, FlowBind learns a straight interpolation path that bridges the data $z^{i}$ to the shared latent $z^{*}$ by:
$z_{t}^{i}$ $=tz^{i}+(1-t)z^{*}$ (4)
$\frac{\partial z^{i}_{t}}{\partial t}$ $=v^{i}(z_{t}^{i},t),$ (5)
where $v^{i}$ denotes the modality-specific velocity field. Note that multi-modal flows are factorized per modality given the shared latent (Eq. 4), and the shared latent implicitly aligns these flows across modalities (Eq. 5). During training, the shared latent is instantiated as $z^{*}=H_{\phi}(\mathbf{z}^{S})$ through an auxiliary encoder $H_{\phi}$, whose marginal approximates $\pi_{\text{shared}}$ and is optimized jointly with the per-modality drift networks $v_{\theta^{i}}$ (Figure 1(a)). At inference, FlowBind relies only on the learned drift networks: owing to the invertibility of the direct flows, both inferring the shared latent from input modalities and generating outputs from the latent are achieved by a single drift network per modality (Figure 1(b)). Details of the training and inference procedures are provided in Section 3.1.
(a) Training both shared latent and drifts.
(b) Inference with per-modality drifts.
Figure 1: An overview of FlowBind. (a) During training, we jointly learn the shared latent and per-modality drift networks in a single stage. (b) At inference, the learned drift networks perform flexible any-to-any generation by solving per-modality ODEs forward and backward in time.
Following prior works esser2024sd3; liu2024audioldm2, FlowBind operates in a compressed latent space obtained by per-modality autoencoders. However, instead of low-level, high-dimensional latent, we adopt compact and semantic representations extracted by strong encoders in each modality, paired with decoders that reconstruct modality-specific details from the encoded feature. This design enables FlowBind to focus on shared structure in a low-dimensional space, making cross-modality alignment simpler and training both faster and efficient.
Taken together, FlowBind provides several advantages over existing approaches. By introducing a shared latent space, FlowBind factorizes the multi-modal flow into independent per-modality drifts, allowing them to operate in isolation with reduced computational cost. This factorization also naturally enables training with arbitrary paired modalities: since each drift network learns only to connect its modality to the shared latent, learning does not depend on specific modalities or fully-paired data. Finally, both the shared latent and modality-specific drifts are optimized jointly with a single flow matching objective, avoiding the multi-stage training pipelines of prior works and yielding a simple and efficient framework.
3.1 Training and Inference
Learning Objective
During training, the auxiliary encoder $H_{\phi}$ and the set of modality-specific drift networks $\{v_{\theta^{i}}\}_{i=1}^{N}$ are optimized jointly under the flow matching framework in Eq. 4 and 5. Given a partially paired sample $\mathbf{z}^{S}$, the auxiliary encoder produces a shared latent $z^{*}=H_{\phi}(\mathbf{z}^{S})$, and for each modality $i\in S$, the drift network $v_{\theta^{i}}$ is trained to approximate the velocity field along the path between $z^{i}$ and $z^{*}$. This leads to the training objective:
$\mathcal{L}(\theta,\phi)=\mathbb{E}_{t,\mathbf{z}^{S},z^{*}}\left[\sum_{i\in S}\left(\left\|v_{\theta^{i}}(z^{i}_{t},t)-(z^{i}-z^{*})\right\|^{2}\right)\right],$ (6)
where $\theta=\{\theta^{1},...,\theta^{N}\}$. In principle, this couples the two components: drift networks learn to predict the displacement toward each modality endpoint, while the auxiliary encoder is encouraged to provide a shared latent from which every modality can be recovered to aid drift networks.
However, this formulation admits degenerate solutions. For example, if the encoder collapses to a constant output such as $z^{*}=0$, the drift networks can trivially fit $v^{i}(z^{i}_{t},t)=z^{i}$ with $z^{i}_{t}=tz^{i}$ and achieve the zero loss for $t\in(0,1]$, leaving the encoder with no meaningful supervision kim2024sfno. The underlying reason is that flow matching enforces transportation between two fixed endpoints but does not itself constrain the distribution of encoder outputs. Prior works on direct flow liu2025crossflow; he2025flowtok address this by adding explicit regularizers, such as contrastive losses on the encoder, but these introduce additional computation, hyperparameters, and scalability bottlenecks especially with increasing number of modalities.
In contrast, we show that both stabilization and meaningful learning of the encoder can be achieved within the flow-matching objective itself. Our approach is simple: for $t\in(0,1]$, we stop gradients through the auxiliary encoder to stably train the drift networks, while at t=0t{=}0, the encoder is directly updated together with the drifts. Despite its simplicity, this scheme effectively prevents collapse and provides the encoder with a meaningful learning signal, as we elaborate below.
Analysis on Encoder Objective
To understand what the auxiliary encoder learns under our training strategy, we analyze the flow matching loss at t=0t{=}0. Substituting the Bayes-optimal drift $v^{\star}(z_{t},t)$ (Eq. 2) into Eq. 6 at $t=0$ gives encoder’s effective objective:
$\mathcal{L}(\phi)=\mathbb{E}\!\left[\|\,v^{\star}(z^{*},0)-(z^{i}-z^{*})\,\|_{2}^{2}\right]=\mathbb{E}\!\left[\|\,\mathbb{E}[z^{i}\mid z^{*}]-z^{i}\,\|_{2}^{2}\right]=\mathbb{E}\!\left[\mathrm{Var}(z^{i}\mid z^{*})\right].$ (7)
This shows that, at t=0t{=}0, the encoder is explicitly optimized to minimize the conditional variance of each modality given the shared latent. The term $\mathbb{E}\!\left[\mathrm{Var}(z^{i}\mid z^{*})\right]$, often referred to as the unexplained variance, measures how much information about modality $i$ remains outside of $z^{*}$. By the law of total variance (grimmett2001lotv), reducing this quantity equivalently increases the explained variance of $z^{i}$ by $z^{*}$. Since the optimization is carried out jointly across all modalities, the encoder is therefore driven to shape $z^{*}$ so that it retains predictive information about each modality, ensuring that the shared latent becomes increasingly informative for cross-modal alignment.
More generally, when the drift networks are not optimal, Eq. 6 at t=0t{=}0 decomposes into unexplained variance and an approximation error of the drifts:
Proposition 1 (Equivalence at $t=0$)
For any parameters $(\theta,\phi)$ and modality subset $S$, the flow matching loss (Eq. 6) at $t=0$ decomposes as follows:
$\mathcal{L}(\theta,\phi)=\underbrace{\sum_{i\in S}\mathbb{E}\!\left[\mathrm{Var}(z^{i}\mid z^{*})\right]}_{\text{unexplained variance}}\;+\;\underbrace{\sum_{i\in S}\mathbb{E}\!\left[\left\|\,v_{\theta^{i}}(z^{*},0)-\mathbb{E}[\,z^{i}-z^{*}\mid z^{*}]\,\right\|^{2}\right]}_{\text{approximation error of }v_{{\theta}^{i}}}.$
A formal proof is provided in Appendix LABEL:app:cond_var_proof.
This decomposition reveals that even when the drift networks are imperfect, the auxiliary encoder $H_{\phi}$ is consistently driven to minimize unexplained conditional variance while simultaneously optimizing the shared latent to enable better drift approximation of their targets. In this way, our training strategy encourages the auxiliary encoder and the drift networks to remain tightly coupled: the drifts learn to predict each modality endpoint from the shared latent ($t\in[0,1]$), while the encoder is driven to shape the latent into a representation from which all modalities can be reliably recovered ($t=0$). During training, we balance the drifts and encoder training by sampling from the mixture $t\sim(1-\alpha)\text{Unif}(0,1)+\alpha\delta(t=0)$. The training procedure is given at Algorithm LABEL:alg:train-no-opt.
Inference
After training, FlowBind performs versatile any-to-any generation relying solely on the learned per-modality flows, without utilizing the auxiliary encoder. Given a source modality $i$, we first project it onto the shared latent by integrating its backward flow, and then map the shared latent to the target modality $j$ via the corresponding forward flow:
$\hat{z}^{*}=\text{ODESolve}(z^{i},v_{\theta^{i}},1,0),\qquad\hat{z}^{j}=\text{ODESolve}(\hat{z}^{*},v_{\theta^{j}},0,1)$ (8)
When conditioning on multiple source modalities $\mathbf{z}^{S}$, FlowBind obtains per-modality latent estimates $\hat{z}^{(*,i)}$ by solving the corresponding backward flows independently. These estimates are then aggregated into the shared latent $\hat{z}^{*}$ by simple averaging. Finally, the target modality is generated by integrating its forward flow starting from $\hat{z}^{*}$. The inference procedure is given at Algorithm LABEL:alg:infer-any2any.
4 Related Work
Any-to-Any Generation
A prominent paradigm for any-to-any generation tokenizes all modalities into a discrete space and trains a single sequence model to predict the unified stream auto-regressively. In this setup, a powerful large language model performs cross-modal sequence generation, with tokenized data of all modalities. Some works (chameleon2024chameleon; orthus2025siqi) focus on interleaved generation solely on text-image generation, while others (wu2024nextgpt; zhan2024anygpt) extend to broader multi-modal scenarios including speech (wang2024mio) and even robotics (lu2024unifiedio2). Training such models typically involves multi-stage procedures and often relies on instruction fine-tuning, which requires datasets with detailed textual descriptions and usually depends on large language models. Consequently, text-paired data is often required for training.
Another line of work utilizes discrete diffusion models, often by adapting them to operate on discrete token spaces (rojas2025diffuse_every; shi2025muddit). These methods, which typically focus on text-image generation tasks, leverage the high-quality synthesis capabilities of diffusion for multi-modal scenarios. For instance, UniDisc (swerdlow2025unidisc) highlights the controllability of this approach by framing various conditional generation tasks, such as inpainting.
Direct Flow-based Models
Recent flow-based models have explored learning direct, data-to-data invertible mappings between two data distributions (li2023bbdm; wang2024semflow), predominantly focusing on text-image pairs (liu2025crossflow; he2025flowtok). This approach represents a fundamental departure from traditional generative flows that typically learn bridging from fixed prior distributions (e.g., standard Gaussian) to target data distributions through conditional generation mechanisms. To facilitate these direct transformations, existing methodologies designate latent distributions of one modality (i.e., source distribution) as a learnable embedding. This is achieved by introducing an encoder for the source modality and constructing additional loss terms that align the source and target modalities, such as contrastive learning objectives.
While our approach shares foundational ideas with prior work, its emphasis and formulation differ. Existing methods typically rely on multiple loss terms to stabilize training and to optimize endpoint embeddings; in contrast, we employ a single, unified flow objective to achieve the same optimization. Moreover, we pursue direct flows for multi-modality connectivity, whereas most prior efforts have concentrated on two-modality settings, especially text–image generation.
5 Experiments
We conduct an extensive evaluations on any-to-any generation tasks across text, image, and audio modalities, covering a wide range of input-ouput modality combinations including one-to-one, one-to-many and many-to-one generation tasks. For baselines, we mainly consider approaches on flow-based any-to-any generative modeling, namely CoDi tang2023codi and OmniFlow li2025omniflow. Qualitative results for various input-output modality combinations are available on our official project page: https://yeonwoo378.github.io/official_flowbind.
Tasks and Evaluation Protocol
For one-to-one generation, we consider all six possible tasks that consist of text, image and audio, and discuss its result in Sec. 5.2. We adopt established evaluation metrics for one-to-one generation, where standard measures of generation quality and cross-modal alignment are available. Specifically, generation quality is assessed using established modality-specific measures: FID (heusel2017fid) for images, FAD (kilgour2019fad) for audio, and CIDEr (vedantam2015cider) for text captions. Cross-modal alignment is evaluated through pairwise similarity metrics: CLIP scores for text-image pairs hessel2021clipscore, CLAP scores for text-audio pairs elizalde2023clap, and Audio-Image-Similarity (AIS) (wu2022wav2clip) for image-audio pairs. Evaluations are done at held-out test set for text-audio and image-audio tasks, while we employ widely adopted zero-shot benchmark in MS-COCO for text-to-image and image-to-text tasks. Detailed evaluation protocols and descriptions of the specialist baseline models are in Appendix LABEL:app:eval.
For many-to-many generation, we conduct a qualitative analysis of more challenging settings to validate FlowBind’s cross-modal capabilities. Furthermore, since no standard benchmark exists for many-to-many generation, we introduce a tailored quantitative evaluation metric and report corresponding results in Sec. 5.3, demonstrating that FlowBind consistently outperforms the baselines.
Implementation Details
We employ EmbeddingGemma (google2025gemma3) for textual semantic latent, CLIP (radford2021clip) for visual latent with Stable-UnCLIP (stableunclip) as decoder, and CLAP (elizalde2023clap) features for audio synthesis conditioning. Note that these modality-specific encoders and decoders are frozen during the training of FlowBind. We employ MLP-based architecture with residual connections for both auxiliary encoders and drift networks, with AdaLN-zero for time modulation peebles2023dit. More detailed information, including the architectural and training specifications, can be found in Appendix LABEL:apx:implementation_detail.
5.1 Instantiation of FlowBind
Table 1: Comparison of computational cost. #(A-B) indicates the number of training samples for each dataset combination. Training time for CoDi is omitted due to absence of training code and details. For OmniFlow, we report the training time only for the final joint training stage.
Model Train Param. GPUhrGPU-hr Number of Traning Data Joint Training
#(T–I) #(T–A) #(I–A) #(T–A–I)
CoDi 4.3B - 400M 3.5M 1.9M - no
OmniFlow 3.2B 480hr480hr* 28M 2.4M - 2.2M no
FlowBind 568M $48hr$ 310K 96K 180K - yes
[…]
5.2 Results on One-to-One Generation
Table 2: Fidelity assessment on one-to-one evaluation benchmarks.
Category  Model  $\mathrm{T}\to\mathrm{I}$ $\mathrm{I}\to\mathrm{T}$  $\mathrm{T}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{T}$  $\mathrm{I}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{I}$
    $\mathrm{FID}\downarrow$ $\mathrm{CIDEr}\uparrow$  $\mathrm{FAD}\downarrow$ $\mathrm{CIDEr}\uparrow$  $\mathrm{FAD}\downarrow$ $\mathrm{FID}\downarrow$
Specialists  SD3-Medium    25.40 –    –
  FLUX.1    22.06 –    –
  LLaVA-NeXT    109.3    –
  TangoFlux    –    1.41 –
  AudioX    –    3.09 –
  Qwen2-Audio    –    4.64
  Seeing & Hearing    –    –    5.31
  Sound2Vision    –    –    42.55
Generalists  UnifiedIO2-L  21.54 134.7*  8.31 12.15
  CoDi    24.80 16.40    9.84 6.62    14.58 50.4
  OmniFlow    22.97 44.20    4.20 31.79    5.67 106.03
[…]
Table 3: Alignment results on one-to-one evaluation benchmarks.
Category  Model  $\mathrm{T}\to\mathrm{I}$ $\mathrm{I}\to\mathrm{T}$  $\mathrm{T}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{T}$  $\mathrm{I}\to\mathrm{A}$ $\mathrm{A}\to\mathrm{I}$
    $\mathrm{CLIP}\uparrow$ $\mathrm{CLIP}\uparrow$  $\mathrm{CLAP}\uparrow$ $\mathrm{CLAP}\uparrow$  $\mathrm{AIS}\uparrow$ $\mathrm{AIS}\uparrow$
Specialists  SD3-Medium    31.60 –    –
  FLUX.1    31.06 –    –
  LLaVA-NeXT    32.14    –
  TangoFlux    –    42.71 –
  AudioX    –    29.29 –
  Qwen2-Audio    –    17.09
  Seeing & Hearing    –    –    75.11
  Sound2Vision    –    –    62.39
Generalists  UnifiedIO2-L  30.71 30.73  13.48 18.68
  CoDi    30.26 26.24    10.79 17.94    61.55 74.26
  OmniFlow    31.52 27.71    24.23 45.08  71.71 59.22
[…]
Effectiveness of FlowBind
We demonstrate the effectiveness of FlowBind under all six pairwise one-to-one generation scenarios in Table 5.2 and Table 5.2. While the core goal of FlowBind lies on efficient modeling of any-to-any generation, we also observe the resulting model shows strong capability in cross-modal generation tasks. Compared to CoDi and OmniFlow, FlowBind achieves the best quality metrics in all six one-to-one generation tasks, while showing superior alignment score on four tasks among six. We also note that baselines such as OmniFlow are initialized from strong specialist model (i.e., SD3-Medium) that excels at text-image alignment, which explains their particularly good performance on text-to-image alignment scores. We compare FlowBind with UnifiedIO2-L (lu2024unifiedio2), a recent LLM-based any-to-any generation method. FlowBind is comparable on text–image tasks and stronger in other settings, suggesting that our formulation may provide benefits beyond flow-based methods. As an overall, we conclude that FlowBind shows promising performance on the evaluated one-to-one generation tasks.
An interesting observation is that FlowBind exhibits substantial gains in the image-audio generation, where it significantly outperforms among generalists and even dedicated specialist, without making modality-specific adjustments. We conjecture the impressive performance of FlowBind at audio-image correspondence stems from the introduction of learnable shared latent space, which is designed to contain meaningful information about each modality (Section 3.1) and learns directly from audio-image pair. Instead of learning a shared latent space from arbitrarily paired data, CoDi employs an text-anchored design, using only text-paired data during its multimodal alignment stage. This design choice of CoDi makes alignment between non-text modality, such as audio-image alignment, to be indirectly captured with the aid of text. We note that OmniFlow is also implicitly relying on text representation, given the fact that its weights are initialized from pretrained text-to-image and text-to-audio models at the beginning of second-stage any-to-any training. In contrast, FlowBind learns a shared latent space directly from the available paired training data, treating all modalities symmetrically and thereby offering a more suitable representation for cross-modal generation.
Train Efficiency
While achieving promising performance relative to prior any-to-any generation models, we emphasize that FlowBind is trained with substantially less computation and effort. As previously shown in Table 1, the demonstrated strong performance of FlowBind is achieved using 6 times less training parameters and 10 times less compute compared to OmniFlow. Moreover, FlowBind employs a unified training objective, in contrast to prior works that require complex multi-stage training pipelines. Consequently, FlowBind can be trained with less effort without cumbersome hyperparameters and additional computations that emerge from those complex training procedures.
Data Efficiency
In terms of data efficiency, FlowBind achieves any-to-any generation with a much smaller training dataset, using 0.15 % of CoDi and 1.79 % of OmniFlow. We conjecture the training can be done with much smaller dataset because we choose to model flow between high-level representations. By doing so, the cross-modal generation capability is decomposed into inter-modal alignment and intra-modal generation in FlowBind. Our drift network is only required to capture inter-modal correspondence, as per-modality frozen encoder-decoders take charge in capturing intra-modal generative capability. This would enable FlowBind to quickly capture cross-modal alignment with smaller datasets.
5.3 Results on Many-to-Many Generation
Figure 2: Qualitative results on various many-to-many generation tasks. More results and comparisons with baselines are presented in Appendix LABEL:app:many-to-many-qual.
Beyond our extensive one-to-one evaluations, we further conduct qualitative analyses to assess capability of FlowBind as an any-to-any generation model on more challenging cross-modal generation tasks, complemented by our own quantitative comparisons against strong baselines. As shown in Figure 2, FlowBind handles these complex settings while faithfully reflecting the input conditions in its generated outputs.
Interestingly, we observe that fine-grained details from the input (e.g., the Stars and Stripes pattern printed on the table) can reappear in the generated modalities, as shown in the second row of Fig. 2(a). This suggests that FlowBind’s learned shared space is expressive enough for cross-modal generation, allowing it to aggregate complementary information from multiple input conditions via latent averaging. We further analyze the effectiveness of this aggregation in Appendix LABEL:app:averaging. Additional qualitative examples are provided in Appendix LABEL:app:many-to-many-qual and on our project page.
We also conduct quantitative evaluations of FlowBind on many-to-one and one-to-many generation tasks to validate its performance and compare it against baselines. To this end, we construct a synthetic triplet dataset by extending the AudioCaps (kim2019audiocaps) text–audio pairs. Following a protocol similar to OmniFlow, we generate the missing image modality using FLUX.1-schnell (BlackForestLabs2024FLUX1), conditioned on the text annotations. This yields a triplet (text, image, audio) dataset that enables quantitative evaluation of many-to-many generation.
[…]
Figure 16: Results on {text+image}-to-image generation.